OpenStudy (31356):

A worm and a snail enter the relay races. If the worm goes for 5 hours and then the snail goes for 4 hours, they cover a total of 448 inches. If the worm goes for 2 hours and the snail goes for 3 hours, they cover a total of 252 inches. How fast does each scoot along? SHOW YOUR WORK! & MEDAL REWARDED!!

OpenStudy (31356):

Anyone?

OpenStudy (anonymous):

you would simply divide 5 hours (for the worm) by 448 to get how much inches is covered per hour

OpenStudy (31356):

I don't think it is that. Since, this is Problem Solving by Systems.

OpenStudy (campbell_st):

why post the question again...

OpenStudy (anonymous):

if you already know how to solve it.. then why did you post the quesiton

OpenStudy (campbell_st):

a lot of people don't give answers... they give advice on how to solve a question... that's the aim of the site, assisting you to improve your understanding

OpenStudy (anonymous):

*applauses* @campbell_st

OpenStudy (radar):

Using the "system" and the process of elimination. Let s = rate of snail in in./hr Let w = rate of worm in in./hr Set up the equations using info provided by the probleml Eq 1: 5s + 4w= 448 in. Eq 2: 3s + 2w = 252 in To eliminate "w" variable X -2 the 2nd Eq. -6s - 4w = -504 now add the 1st Eq to eliminate the w term 5s +4w = 448 ---------------- -s = -56 s=56 Snail rate is 56 in. per hr. Now substitute this value for snail in original equation and solve for worm rate.